Hi
jack and jill were each given the same amount of money. Every day jack spent 24 dollars and jill spent 20dollars.when jack ran out of money, jill still had 36 dollars left. How much was given to each
initially.
thanks

Answer by Theo(13342)

let y = the starting balance.
let x = the number of days.

you get two equations.

y - 24x = 0
y - 20x = 36

subtract the first equation from the second to get:

4x = 36

divide both sides of this equation by 4 to get x = 9

solve each equation after replacing x with 9 to see what you get.

y - 24x = 0 becomes y - 24*9 = 0 which becomes y - 216 = 0 which becomes y = 216.

y - 20x = 36 becomes y - 20*9 = 36 which becomes y - 180 = 36 which becomes y = 216.

your starting balance is 216.
the number of days is 9.

216 - 24*9 = 0
216 - 20*9 = 36